OpenStudy (anonymous):

find the roots of the quadratic equation y = 3x2 - 30x + 45 .

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

remember we have to help her not give her the answer

OpenStudy (anonymous):

let's take 3 as common factor so what we get y=3(x2-10x+15) now we will find roots of equation by formula method x=-b+-sqrt(b^2-4ac)/2a .................(1) by comparing our quadratic equation x2-10x+15 to standard form ax2+bx+c we get a=1 b=-10 c=15 now put this values in (1) and get values of x and put them in to our given equation y=3(x2-10x+15)

OpenStudy (anonymous):

thanks wolf923

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

you are most welcome